# Principal stratification with $U$-statistics under principal ignorability

### v4

@@ 6.1 SPI for continuous outcomes
We first implement the proposed methods by considering the job status one year after the assignment as the intermediate outcome, $D$, where if participant $i$ is employed one year after the assignment, then $D_{i}=1$ and the final outcome $Y_{i}$, the weekly earning, can be observed and defined without ambiguity; otherwise, $D_{i}=0$ and $Y_{i}$ is not well-defined and denoted by $Y_{i}=*$. In other words, we focus on the initial random assignment (an intention-to-treat analysis) and address earnings outcomes truncation by “death” (employment status); also see Zhang et al. (2009) for a similar setup. Under this setting, we estimate the SPI estimand defined in (3) for the always-employed (principal stratum $11$), i.e., participants who will be employed with or without Jobs Corps. For methods based on parametric working models, a logistic model is considered for the employment process within each assignment group, and a log-normal model is fitted for the final earning outcome to induce the pairwise mean outcome function. For the debiased machine learning approach, we use a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner with the random forest and generalized linear models libraries.
The estimated results for SPI are presented in Table 2, where $1,000$ bootstrap replicates were generated to estimate the standard errors. We observe that, for the always-employed (72.9% of all participants by estimation), the probabilistic index is estimated to be $0.539$ using the triply robust and debiased machine learning methods. This indicates that the probability that a randomly chosen always-employed under assigned treatment has a higher-ranked earning outcome than a randomly chosen always-employed under control is $0.539$, with the 95% confidence interval excluding the null of $0.5$; this implies that the Jobs Corps positively affects one’s earnings in the one year after the assignment. Finally, the point estimates obtained under the plug-in moment methods are slightly closer to the null compared to those from the triply robust and debiased machine learning methods, although their 95% confidence intervals continue to exclude the null.
@@ 6.2 Principal win ratio (PWR) for ordinal outcomes
We provide additional analysis of the Job Corps data to illustrate the application of the proposed methods to address ordinal outcomes with noncompliance. We use the actual treatment status as an intermediate variable, similar to Chen and Flores (2015). Under the monotonicity assumption, three strata potentially exist, including the compliers (stratum $10$), the never-takers (stratum $00$), and the always-takers (stratum $11$). To address the possibility of unemployment status at the time of earning assessment, we create an ordinal outcome with three ordered categories: 0 (unemployment), 0 to 170 per week (hourly wage 4.25 dollars on a 40-hour-per-week basis), and above 170 dollars per week, following the categorization considered by Lu et al. (2018). Under this setting, we estimate the principal win ratios within each principal stratum, as defined in (4). For methods based on parametric working models, the pairwise outcome mean is estimated using ordered logit regression, while treatment propensity and principal scores are estimated using logistic regression. For the debiased machine learning approach, we used multinomial generalized linear models with the elastic net penalty to estimate the pairwise outcome mean. The treatment propensity and principal scores were estimated using the SuperLearner, fitted under a five-fold cross-fitting ($K=5$ and $L=15$) procedure.
@@
The estimated results are presented in Table 3, where $1,000$ bootstrap replicates were generated to estimate the standard errors. The estimated proportions of always-takers, compliers, and never-takers among all participants are 50.1%, 34.0%, and 15.9%, respectively. The Jobs Corps intervention assignment has a positive effect on participants’ earnings one year after the assignment for both compliers and always-takers, and this effect appears to be more pronounced for compliers. For example, using the debiased machine learning estimator, there is a 27% increase in the chances of being in a higher earning category for a complier under treatment compared to a complier under control. The 95% lower confidence limits obtained from all five methods are strictly above the null value of 1. For the never-takers, however, the effect of Jobs Corps appears more neutral, as the principal win ratio is estimated to be closer to 1, and three out of five methods (M1, triply robust and debiased machine learning) produce 95% confidence intervals that include the null value of 1. Interestingly, in this reanalysis using a win ratio estimand, we find that treatment effect heterogeneity may exist across the three compliance strata, and the exclusion restriction assumption often invoked for compliance analysis may be questionable. The statistically significant effect among the always-takers is indicative of a positive psychological effect. That is, participants assigned to the Jobs Corp intervention may feel encouraged and naturally be more optimistic about future employment opportunities (hence more likely to move into a higher earning category) compared to those who switched from their initial control assignment.
@@ 6.3 Sensitivity analyses for monotonicity
We conducted additional sensitivity analyses to examine the impact of potential violations of the monotonicity assumption, focusing on both Sections 6.1 and 6.2. Following the approach in Section 5.3, we specified the sensitivity function as $\eta(X)=\eta_{0}[1-\{p_{1}(X)-p_{0}(X)\}/\min\{p_{1}(X),1-p_{0}(X)\}]$, with $\eta_{0}\in\{0.1,0.2,0.3,0.4,0.5\}$. This formulation allows the sensitivity function to naturally vary with covariates while being governed by a one-dimensional strength parameter $\eta_{0}$. The chosen range of $\eta_{0}$ represents mild to moderate departures from monotonicity, which is plausible in the context of the Jobs Corps study, where encouragement for employment and subsequent treatment receipt may not dramatically deviate from monotonicity. We applied both the triply robust and the debiased machine learning estimators, with results reported in Figure 3.
For the SPI analysis (Figure 3(a)), increasing $\eta_{0}$—corresponding to stronger departures from monotonicity—led to slightly larger estimated PGCEs. In other words, as the proportion of harmed strata increases, the estimated SPI shifts further from the null, yielding a stronger treatment effect signal, though statistical significance remains unaffected. In contrast, for the PWR analyses (Figures 3(b) - (d)), the estimated PGCEs are largely stable across values of $\eta_{0}$. Despite the presence of defiers, the estimated win ratios across all three principal strata remain nearly unchanged. The only noticeable pattern is a narrowing of the 95% confidence interval for the complier win ratio as $\eta_{0}$ increases. Collectively, these findings suggest that the PWR estimators exhibit strong robustness to non-monotonicity, in contrast to the slightly more pronounced sensitivity observed for the SPI.
(a) SPI
(b) $\mathrm{PWR}^{10}$
(c) $\mathrm{PWR}^{00}$
(d) $\mathrm{PWR}^{11}$
Figure 3: Estimated SPI and PWRs and 95% confidence intervals under different values of sensitivity analysis parameter $\eta_{0}$. The horizontal axis of each plot shows values of $\eta_{0}$, and the vertical axis shows values of the estimated estimands and confidence intervals. SPI: the survivor probabilistic index in Section 6.1; $\mathrm{PWR}^{s}$: the principal win ratio of stratum $s$ in Section 6.2, with $s\in\{10,00,11\}$.
@@ 7 Discussion
In this article, we develop principal stratification methods for addressing a general class of PGCE estimands defined in terms of a binary intermediate variable. Under the assumptions of treatment unconfoundedness, monotonicity, and mean principal ignorability that are based on the contrast function, we contributed a semiparametric efficient approach for estimating the PGCE estimands by identifying their efficient influence functions and obtaining the ratio $U$-statistic estimators that combine several different nuisance working functions. We showed that the proposed $U$-statistic estimators are triply robust and become locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds when all nuisance functions are correctly specified by parametric models or machine learning models with cross-fitting.
@@
As in all principal stratification analyses that achieve point identification, the proposed $U$-statistic estimators for estimating PGCEs also depend on the mean principal ignorability assumptions that are unverifiable from the observed data alone. In Section S7 of the Online Supplement, we have described an additional sensitivity function approach to obtain the bias-corrected $U$-statistic estimators for estimating the PGCEs when the mean principal ignorability assumption is violated. The bias-corrected estimators remain triply robust to potential working model misspecifications and provide means to generate a range of point estimates under assumed non-ignorability.
@@
Finally, our current development assumes that the observed data vectors are independent and identically distributed, providing a natural starting point. Extending these ideas to correlated data structures is an important direction for future research. For clustered data, recent work in cluster randomized trials has emphasized the need to define estimands that explicitly acknowledge the multilevel structure, particularly in the presence of informative cluster size (Kahan et al., 2023). A promising avenue is to expand the semiparametric and nonparametric methods of Wang et al. (2024) to accommodate cluster-average and individual-average PGCE estimands when the intermediate variable is measured at the individual level. For longitudinal data, both intermediate and final outcomes may be repeatedly measured over time, which requires reformulating the identification assumptions to incorporate sequential ignorability of treatment and multilevel monotonicity (Frangakis et al., 2004). It is also worth noting that, while our main exposition has focused on a single final outcome, our framework naturally extends to multiple hierarchically ordered outcomes—a setting that has gained considerable traction in clinical research over the past decade. Hierarchical composite endpoints can be viewed as clustered outcome vectors within individuals. In Section S8 of the Online Supplement, we extend the notation of Bebu and Lachin (2015) to the potential outcomes framework and show how our results apply to this setting by defining pairwise comparison estimands with hierarchical outcome components while maintaining the same binary intermediate variable. Further methodological developments along these directions will continue to broaden the applicability of our framework.

### v5

@@ 6.1 SPI for continuous outcomes
We first implement the proposed methods by considering the job status one year after the assignment as the intermediate outcome, $D$, where if participant $i$ is employed one year after the assignment, then $D_{i}=1$ and the final outcome $Y_{i}$, the weekly earning, can be observed and defined without ambiguity; otherwise, $D_{i}=0$ and $Y_{i}$ is not well-defined and denoted by $Y_{i}=*$. In other words, we focus on the initial random assignment (an intention-to-treat analysis) and address earnings outcomes truncation by “death” (employment status); see also Zhang2008 for a similar setup. Under this setting, we estimate the SPI estimand defined in (3) for the always-employed (principal stratum $11$), i.e., participants who will be employed with or without Jobs Corps. For methods based on parametric working models, a logistic model is fitted for the employment process within each assignment group, and a log-normal model is fitted for the final earnings outcome to estimate the pairwise mean outcome function. For the debiased machine learning approach, we use a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner with the random forest and generalized linear models libraries.
The estimated results for SPI are presented in Table 2, where $1,000$ bootstrap replicates were generated to estimate the standard errors. We observe that, for the always-employed (72.9% of all participants, by estimation), the probabilistic index is estimated at $0.539$ using the triply robust and debiased machine learning methods. This indicates that the probability that a randomly chosen always-employed under assigned treatment has a higher-ranked earning outcome than a randomly chosen always-employed under control is $0.539$, with the 95% confidence interval excluding the null of $0.5$; this shows that the Jobs Corps positively affects one’s earnings in the one year after the assignment. Finally, the point estimates obtained under the plug-in moment methods are slightly closer to the null compared to those from the triply robust and debiased machine learning methods, although their 95% confidence intervals continue to exclude the null.
@@ 6.2 Principal win ratio (PWR) for ordinal outcomes
We provide additional analysis of the Job Corps data to illustrate the application of the proposed methods to address ordinal outcomes with noncompliance. We use the actual treatment status as an intermediate variable, similar to Chen2015. Under the monotonicity assumption, three strata potentially exist, including the compliers (stratum $10$), the never-takers (stratum $00$), and the always-takers (stratum $11$). To address the possibility of unemployment status at the time of earning assessment, we create an ordinal outcome with three ordered categories: 0 (unemployment), 0 to 170 per week (hourly wage 4.25 dollars on a 40-hour-per-week basis), and above 170 dollars per week, following the categorization considered by Lu2018. Under this setting, we estimate the principal win ratios within each principal stratum, as defined in (4). For methods based on parametric working models, the pairwise outcome mean is estimated using ordered logit regression, while treatment propensity and principal scores are estimated using logistic regression. For the debiased machine learning approach, we used multinomial generalized linear models with an elastic net penalty to estimate the pairwise outcome mean. The treatment propensity and principal scores were estimated using the SuperLearner, fitted via a five-fold cross-fitting ($K=5$ and $L=15$) procedure.
@@
The estimated results are presented in Table 3, where $1,000$ bootstrap replicates were generated to estimate the standard errors. The estimated proportions of always-takers, compliers, and never-takers among all participants are 50.1%, 34.0%, and 15.9%, respectively. The Jobs Corps intervention assignment has a positive effect on participants’ earnings one year after the assignment for both compliers and always-takers, and this effect appears to be more pronounced for compliers. For example, using the debiased machine learning estimator, there is a 27% increase in the probability of being in a higher-earning category among compliers under treatment relative to those under control. The 95% lower confidence limits obtained from all five methods are strictly above the null value of 1. For the never-takers, however, the effect of Jobs Corps appears more neutral, as the principal win ratio is estimated to be closer to 1, and three out of five methods (M1, triply robust and debiased machine learning) produce 95% confidence intervals that include the null value of 1. Interestingly, in this reanalysis using a win ratio estimand, we find that treatment effect heterogeneity may exist across the three compliance strata, and the exclusion restriction assumption often invoked for compliance analysis may be questionable. The statistically significant effect among the always-takers indicates a positive psychological effect due to the assignment alone. That is, participants assigned to the Jobs Corp intervention may feel encouraged and naturally be more optimistic about future employment opportunities (hence more likely to move into a higher-earning category) than those who switched from their initial control assignment.
@@ 6.3 Sensitivity analyses for monotonicity
We conducted additional sensitivity analyses to examine the impact of potential violations of the monotonicity assumption, focusing on both Sections 6.1 and 6.2. Following the approach in Section 5.3, we specified the sensitivity function as $\eta(\bm{X})=\eta_{0}[1-\{p_{1}(\bm{X})-p_{0}(\bm{X})\}/\min\{p_{1}(\bm{X}),1-p_{0}(\bm{X})\}]$, with $\eta_{0}\in\{0.1,0.2,0.3,0.4,0.5\}$. This formulation allows the sensitivity function to naturally vary with covariates while being governed by a one-dimensional strength parameter $\eta_{0}$. The chosen range of $\eta_{0}$ represents mild to moderate departures from monotonicity, which is plausible in the context of the Jobs Corps study, where encouragement for employment and subsequent treatment receipt may not deviate dramatically from monotonicity. We applied both the triply robust and the debiased machine learning estimators, with results reported in Web Figure S5 in Section S7.2 of the Online Supplement.
For the SPI analysis (Web Figure S5(a)), increasing $\eta_{0}$, corresponding to stronger departures from monotonicity, led to slightly larger estimated PGCEs. In other words, as the proportion of harmed strata increases, the estimated SPI shifts further from the null, yielding a stronger treatment effect signal, although statistical significance remains unchanged. In contrast, for the PWR analyses (Web Figures S5(b) - (d)), the estimated PGCEs are largely stable across values of $\eta_{0}$. Despite the presence of defiers, the estimated win ratios across all three principal strata remain nearly unchanged. The only noticeable pattern is a narrowing of the 95% confidence interval for the complier win ratio as $\eta_{0}$ increases. These findings suggest that the PWR estimators exhibit strong robustness to non-monotonicity, in contrast to the slightly more pronounced sensitivity observed for the SPI.
@@ 7 Discussion
We develop principal stratification methods for addressing a general class of PGCE estimands defined in terms of a binary intermediate variable. Under the assumptions of treatment unconfoundedness, monotonicity, and mean principal ignorability based on the contrast function, we proposed a semiparametric approach for estimating the PGCEs by identifying their efficient influence functions and obtaining scaled U-statistic-based estimators. We showed that the proposed U-statistic-based estimators are triply robust and locally efficient, with asymptotic variances that attain the semiparametric efficiency lower bounds when all nuisance functions are correctly specified by parametric or machine learning models with cross-fitting.
@@
As in all principal stratification analyses that achieve point identification, the proposed U-statistic-based estimators of PGCEs also depend on the mean principal ignorability assumptions, which are unverifiable from the observed data alone. In Section S7 of the Online Supplement, we describe an additional sensitivity-function approach to obtain bias-corrected U-statistic-based estimators of the PGCEs when the mean principal ignorability assumption is violated. The bias-corrected estimators remain triply robust to potential working model misspecifications and provide a means to generate a range of point estimates under the assumption of non-ignorability.
@@
Finally, our current development assumes that the observed data vectors are independent and identically distributed, providing a natural starting point. Extending these ideas to correlated data structures is an important direction for future research. For clustered data, recent work on cluster randomized trials has emphasized the need to define estimands that explicitly account for the multilevel structure, particularly in the presence of informative cluster size (kahan2023estimands). A promising avenue is to expand the semiparametric and nonparametric methods of wang2024model to accommodate cluster-average and individual-average PGCE estimands when the intermediate variable is measured at the individual level. For longitudinal data, both intermediate and final outcomes may be repeatedly measured over time, which requires reformulating the identification assumptions to incorporate sequential ignorability of treatment and multilevel monotonicity (frangakis2004methodology). It is also worth noting that, while our main exposition has focused on a single final outcome, our framework naturally extends to multiple hierarchically ordered outcomes, a setting that has gained considerable traction in clinical research over the past decade. In Section S8 of the Online Supplement, we extend the notation of Bebu2015 to the potential outcomes framework and show how our results apply to this setting by defining pairwise comparison estimands with hierarchical outcome components while maintaining the same binary intermediate variable. Further methodological developments along these directions will continue to broaden the applicability of our framework.